# TESSERACT Collaboration

### v2

@@ I Motivation
At present, the field lacks an understanding of the fundamental processes setting the background environment and energy resolution of such detectors. For nearly the past decade, low threshold phonon detectors have observed orders of magnitude more events below several hundred eV than expected from background radioactivity or cosmic ray interactions Adari et al. (2022); Baxter et al. (2025), events often called the “low energy excess” (or LEE). These excess events couple both directly to a detector’s sensor films (“singles” LEE), and more problematically, to a detector’s phonon system (“shared” LEE) Anthony-Petersen et al. (2025); Angloher et al. (2024a). Similarly, our collaboration has observed excess noise in our phonon sensors, consistent with sub-threshold events in the phonon system (“correlated” or “phonon” noise) and our phonon sensors themselves (“uncorrelated” noise) Anthony-Petersen et al. (2025).
Superconducting qubits and similar devices have observed quasiparticle bursts in their superconducting films Mannila et al. (2022) including correlated bursts seemingly caused by phonon bursts in the substrateYelton et al. (2025). While potential connections to LEE-like phenomena have been proposedBaxter et al. (2025), the precise link between observations in qubits and low threshold particle detectors remains unclear.
A variety of explanations have been proposed for observations in low threshold phonon detectors, including the relaxation of thermal stress in the metal films used to construct phonon sensorsRomani (2024), or defects in the detector’s bulk crystalNordlund et al. (2024) or on its surface. This work probes the scaling of these backgrounds and noise terms with the detector volume (or equivalently mass).
@@ II Experimental Setup
Figure 1: (Top left) Photograph of our 1 mm (left) and 4 mm thick (right) 1 cm2 silicon detectors. (Top right) Detail on mask design for our phonon sensors (QETs). For scale, the QET fins have a radius of approximately 140 $\mu$m. (Bottom) Sketch of backgrounds and shot noise sources we observe in our devices. Shared LEE backgrounds and correlated phonon shot noise appear at approximately 4x the rate in the 4 mm detector compared to the 1 mm detector.
To calibrate our detectors’ response, we illuminate them from above with pulses of 450 nm (2.755 eV) photons, diffused from the tip of a fiber which runs from the detector cavity to a room temperature laser. We trigger calibration events on a logic level signal that rises simultaneously with the firing of the laser, and analyze the response in each channel using optimum filtersAnthony-Petersen et al. (2025); Chang et al. (2025). Plotting the height of the resulting events in each channel (Fig. 2 left) , we observe the same features as in Ref. Anthony-Petersen et al. (2025): most photons are absorbed in the detector phonon system, and produce approximately equal responses in both channels, while occasionally photons are directly absorbed into the aluminum phonon collection fins, producing a large response only in one channel. Because multiple photons may strike the detector in one laser pulse, we also observe a superposition of these event types.
To combine the responses of the two channels of one detector, we use a “2x1” multi-channel optimum filter which simultaneously fits the response in both channels, scaling them together by one amplitude Anthony-Petersen et al. (2025); Chang et al. (2025). In our final dataset, we obtain a baseline phonon energy resolution of 258.5$\pm$0.4 meV for this combined channel response, improving on the resolution of our previously world leading detectorAnthony-Petersen et al. (2025); Chang et al. (2025) and previous work by other groupsMichielin (2024); Angloher et al. (2024b).
We periodically calibrate our detector, and find that the energy resolution of both detectors monotonically improves over the course of our run (Fig. 2 Bottom Right). In principle, this could be due to two factors: an improvement in the sensors phonon collection efficiency over time, or an improvement of noise in the detector. Monitoring our detectors’ phonon collection efficiencies over time, we find that they are approximately constant (see supplementary material section A), implying that our noise environment is improving with time.
@@ III Measurement and Modeling of Detector Noise
To characterize our detectors’ noise over time, we take 2 hours of continuous data following each calibration dataset, and remove periods with abnormally small bias power (from periods with abnormally high environmental vibration or after large energy depositions). Additionally, to select periods without above threshold events that would contaminate and bias our noise, we find the largest magnitude pulse in the entire 200 ms period using an optimal filter, and select periods for which this largest fit magnitude pulse has a negative amplitude, i.e. the largest event in the trace is a statistical fluctuation rather than a real energy deposition. In total, 3.9% of the 200ms traces pass these very strict noise selection criteria and are used to estimate the noise Cross power Spectral Density (CSD, see supplementary material section C for further discussion of our cuts). We convert this CSD into the power domain by applying a three-pole responsivity modelLindeman et al. (2011); Irwin and Hilton (2005) for each channel (i.e. $\partial P/\partial I(f)$) that is estimated from IV and $\frac{\partial I}{\partial V}$ measurements as in Ref. Anthony-Petersen et al. (2025).
Similar to our observations in Ref. Anthony-Petersen et al. (2025), we observe that our noise is significantly in excess of modeled TES noise, both in the individual channel Power Spectral Densities (PSDs), and in noise which is correlated between the two channels of the detector (which should be statistically consistent with zero). The broad spectrum correlated noise is consistent with shot noise from many sub-threshold phonon events in the substrate, i.e. it has the same shape and energy partition (50% left, 50% right) as substrate-interacting calibration photons (Fig. 3 Top Right). The magnitude of this broad correlated noise appears to scale with the detector volume, i.e. the correlated noise is four times larger in the 4 mm detector than in the 1 mm detector (Fig. 3 Top Right). The correlated noise peaks in the 100Hz-1kHz range are harmonics of the vibrational resonant modes of the hanging wire bond-detector system.
@@
Figure 3: (Left) Noise in the left channels of the 1 mm (top, green) and 4 mm (bottom, blue) detectors. Our noise is well in excess of our modeled TES noise (gray, dotted), and is composed of correlated phonon noise and uncorrelated noise as in Ref. Anthony-Petersen et al. (2025). The peaks around 150 Hz are due to vibration coupled noise. (Top Right) Correlated noise (i.e. off diagonal CSD element $S_{LR}$) in the 4 mm and 1 mm detectors. The correlated noise in the 4 mm detector is 4 times as large as the 1 mm detector. The amplitude of the 1 mm correlated noise is fit in the orange highlighted region (to avoid vibration coupled peaks), while the 4 mm noise is fit in the orange and yellow regions. (Bottom Right) Uncorrelated noise (i.e. $S_{LL}$, with the modeled correlated noise subtracted) in the left channels of the 1 mm and 4 mm detectors, consistent with the modeled TES Johnson noise at high frequencies and an excess noise term that is approximately flat and consistent between the two channels at low frequencies (see section D for further discussion).
@@
We additionally find that the bias power (i.e. the power applied to the TES to keep it in transition) of our TESs increases over time (see Fig. 4). We assume that this increase over time is essentially entirely due to a shifting parasitic power over time, as shifts in bias power due to a cooling dilution fridge base (mixing chamber, MC) temperature should be very small (see supplementary material section B). As with the correlated noise, the shift in the bias power over time is approximately four times larger in the 4 mm detector than in the 1 mm detector.
The observed monotonic decreases in both the excess noise in the TES and the parasitic power incident on the TES suggest that one source or a set of related sources might cause both phenomena. As the simplest benchmark model, quanta of energy $\varepsilon$ occurring at Poisson-distributed times with an average rate $R(t)\propto t^{-\kappa}$ would induce parasitic bias powers and excess noise terms
@@
where we assume the rate of events $R(t)$ follows a power law, and the rate scales with the detector thickness. (An exponential trend is not a good fit to the data.)
Figure 4: Correlated noise level (Top Left) and Left + Right channel bias powers (Bottom Left) in the 1 mm (green) and 4 mm (blue) detectors as a function of time. (Right) Correlated noise level vs. Left + Right bias powers. Dashed lines show fit models, representing shot noise that scales in rate with detector volume with a shot noise quantum of $\varepsilon=0.68\pm 0.38$ meV.
We plot the excess correlated noise and bias power as a function of time and each other, and find good agreement with this model, which is fit simultaneously to the 1 mm and 4 mm detector data assuming a factor of 4 rate difference. We find $\varepsilon=0.68\pm 0.38$ meV, and $\kappa=0.635\pm 0.009$. Certainly, there is no reason to assume that the shot noise events all have an identical energy. As discussed in supplementary material section E, $\epsilon$ is readily generalizable to processes which have a distribution of energy depositions, where $\varepsilon=\frac{<E^{2}>}{<E>}$. So long as the distribution is not e.g. double peaked, $\varepsilon$ can still be thought of as a characteristic energy scale of the process.
Figure 5: (Top) The mass normalized shared background rates in the 1 mm and 4 mm detectors, with backgrounds measured in the SPICE 1%Anthony-Petersen et al. (2025); Chang et al. (2025), CPDv1Fink et al. (2021); Alkhatib et al. (2021), CPDv2 (previously unpublished), and CRESST 0.35g Si Angloher et al. (2023) detectors. Broadly, all six detectors seem to observe the same backrounds normalizing for mass. Note that CPD and CRESST detectors are one channel detectors which cannot reject singlesAnthony-Petersen et al. (2025), the increased background rates near threshold could be due to singles, as well as noise triggers or additional LEE backgrounds. (Top insert) Detail of the 1 mm, 4 mm, and SPICE 1% detector backgrounds. The orange and green bands show the energy ranges in which time dependence was measured (see bottom right). (Left Bottom) Background event energy partitioning between left and right channels, showing single and shared events. Note the energy reconstruction assumes a shared pulse shape in each channel. (Right bottom) Rate of shared events over time in the 1 mm detector in the shaded 2.5-5 eV and 5-10 eV bins in the center top figure. Dashed line shows the (weighted) average rate in the 5-10 eV range, and a power law fit in the 2.5-5 eV bin.
@@ IV Above Threshold Shared Backgrounds
In addition to this phonon shot noise from below threshold events, we observe a high rate of above threshold background events (often called the “Low Energy Excess,” or LEEAdari et al. (2022); Baxter et al. (2025)). To investigate these LEE events, we recorded 2 or 12 hour periods of continuous background data following each calibration run, and triggered this dataset offline using an optimum filter optimized to search for events with a channel energy partition and pulse shape consistent with that of substrate absorbed calibration photons as in Ref. Chang et al. (2025). In Fig. 5 bottom left, a scatter plot shows the energy deposited in each channel without a constraint on the partition. Identical to observations in Ref. Anthony-Petersen et al. (2025), the LEE backgrounds split into 2 classes: “shared” events with a equal channel partition and a pulse shape identical to substrate absorbed photons, and “singles” events where nearly all energy is deposited in a single phonon sensor channel. Singles have a pulse shape with a fast rise, but a slow fall consistent with an energy deposition that saturates a subset of the athermal phonon sensors that are read out in parallel within a channel. Each event is then fit with a multi-channel optimum filter whose pulse shape and channel partition are fixed to the average of each LEE class (singles left, singles right, shared), and classified according to lowest $\Delta\chi^{2}$.
We see that the rate of above-threshold shared background events scales with the detector mass, as it does for below threshold phonon shot noise (the y-axis has been normalized by substrate mass). However, the rate of these above threshold shared events appears to be constant with respect to time at high energies. At lower energies, where statistics are better, the rate of shared LEE events does appear to decrease over time.
As in Ref. Anthony-Petersen et al. (2025), our singles appear to originate within the aluminum QET films, and we additionally observe that their rate decreases with time (see further discussion in supplementary materials section F). These observations give additional support to the proposal that the dislocation mediated relaxation of thermal stress in these aluminum QET films is responsible for these singles. Anthony-Petersen et al. (2025); Romani (2024)
@@ V Discussion
Our observation of above threshold backgrounds and phonon shot noise which scales with the mass of our detector substrate leads us to conclude that these phonon bursts originate within the detector substrate itself. Through comparisons to detectors with different masses, surface-to-volume ratios, and holding methods,Anthony-Petersen et al. (2025); Chang et al. (2025); Alkhatib et al. (2021); Angloher et al. (2023) we can reject possible LEE scalings with the sidewall area, total surface area, or the holding method, either above (see Fig. 5) or below threshold (see supplementary materials section H). We also conclusively reject the hypothesis that our aluminum sensor films constitute our dominant source of phonon bursts,Romani (2024) but note that such a source could contribute at a subdominant level.
@@
Such a mass scaling suggests that defects in the detector substrate are responsible for LEE and excess phonon noise observations, and is compatible with the model proposed in Ref. Nordlund et al. (2024). The strong rate vs. time since cooldown dependence we observe in our noise data is easily explained by assuming that the relevant excited states are thermally populated at 300K ( >> 1 meV), and then tunnel to lower energy states at mK temperatures. The observation of time dependence in above threshold LEE rates, both here and previouslyAngloher et al. (2023), is more difficult to explain given their eV energy scales are much higher than 300K $k_{B}$. A more complex model, perhaps involving the avalanche-like relaxation of many lower energy states, is needed to explain this time dependence in above threshold data.
The low energy phonon bursts we observe may also constitute the quasiparticle bursts and “poisoning” observed in superconducting qubits and similar devices. We can estimate the reduced quasiparticle density $x_{qp}=n_{qp}/n_{cp}=n_{qp}/(4\times 10^{6}\mu\mathrm{m}^{-3})$Martinis (2021) created by such phonon bursts by assuming that a constant power density $\rho_{p}$ of phonon events originating in the substrate is evenly divided over a $\sim$ 100 nm thick ground and qubit layer which covers essentially all of one face of a superconducting qubit chip. We assume that quasiparticle dynamics in the qubit and ground plane are dominated by recombination (rather than trapping), and that the system is in steady state, such that
$\frac{dx_{qp}}{dt}=0=-rx_{qp}^{2}+g$ (5)
@@
$x_{qp}=\sqrt{\frac{g}{r}}=\sqrt{\frac{\rho_{p}}{\Delta_{Al}n_{cp}}\frac{t_{chip}}{t_{Al}}\frac{1}{r}}\approx 7\times 10^{-8}$ (6)
which is similar to the reduced quasiparticle density observed by well shielded silicon devices Mannila et al. (2022). Our observations also agree with the phonon origin Yelton et al. (2025) and reduction in rate over time Mannila et al. (2022) of such bursts.

### v3

@@ I Motivation
Unfortunately, today’s low-threshold phonon detectors observe orders of magnitude more background events below several hundred eV than expected from background radioactivity or cosmic ray interactions Adari et al. (2022); Baxter et al. (2025). These events are often referred to as the “low energy excess” (LEE). Measurements Angloher et al. (2024a); Anthony-Petersen et al. (2025) show that the LEE can be categorized into two distinct subclasses based on the channel energy partition in detectors with multi-channel phonon sensor readout. “Single” LEE events are found to deposit nearly all of their energy in a single phonon sensor channel, suggesting that they originate within the metal films of the phonon sensors themselves. By contrast, “shared” LEE events deposit nearly equal amounts of energy in all phonon sensor channels, which suggests a substrate origin. Similarly, our collaboration has observed both channel-correlated and uncorrelated excess noise in our phonon sensors, consistent with shot noise from sub-threshold events from these two classes Anthony-Petersen et al. (2025). This work probes the scaling of these backgrounds and noise terms with substrate thickness to further understand their origin.
@@ II Experimental Setup
Here, we measure background events and noise in two nominally identical 1 cm2 silicon low-threshold athermal phonon detectors, where the silicon substrate thickness was varied from 1 mm (mass: 0.233 grams) to 4 mm (mass: 0.932 grams). To read out phonon signals, the detectors use tungsten Transition Edge Sensors (TES, $T_{c}\approx$ 50 mK) coupled to aluminum phonon collection fins in the common Quasiparticle-trap-enhanced Electrothermal-feedback Transition Edge Sensor (QET) Irwin et al. (1995) architecture, employing an identical design to the detectors in Ref. Anthony-Petersen et al. (2025). These individual phonon sensors are aggregated into two separately read out channels to distinguish events and noise which deposit energy in a single phonon sensor from those that deposit energy in multiple phonon sensors Anthony-Petersen et al. (2025); Angloher et al. (2024a); Romani (2023). The phonon sensors on these detectors have improved phonon energy collection efficiency, saturation energy, and channel uniformity compared to those from an earlier fabrication run that suffered from accidental tungsten over-etching Anthony-Petersen et al. (2025). As in Ref. Anthony-Petersen et al. (2024), both detectors are suspended by wire bonds to suppress backgrounds associated with detector holding and housed together in an IR and EMI shielded set of housings attached to the mixing chamber stage of a dilution refrigerator. They were read out using single-stage DC SQUID array amplifiers.
@@
[…]
To calibrate our detectors’ response, we illuminate the phonon sensor face with pulses of 450 nm (2.755 eV) photons, emitted from a diffuser at the tip of a single mode fiber which runs from the detector cavity to a room temperature laser. To precisely record the time of photon calibration events, a logic level signal from the signal generator that powers the laser is concurrently recorded and stored with the continuously recorded signal data streams. Phonon signal amplitudes are then estimated offline using optimum filtersAnthony-Petersen et al. (2025); Chang et al. (2025). Plotting the height of the resulting events in each channel (Fig. 2 left), we observe the same features as in Ref. Anthony-Petersen et al. (2025). Most photons are absorbed in the detector substrate system and produce approximately equal responses in both channels, while occasionally photons are directly absorbed by the aluminum phonon collection fins, producing a large saturated response in only one channel. Because multiple photons may strike the detector in one laser pulse, we also observe a superposition of these event types.
[…]
We periodically calibrate our detector and find that the energy resolution of both detectors monotonically improves over the course of the run (Fig. 2 Bottom Right). In principle, this could be due to two factors: an improvement in the sensors’ phonon collection efficiency over time, or a reduction in noise in the detector. Monitoring our detectors’ phonon collection efficiencies over time, we find that they are approximately constant (see supplementary material section A), implying that our noise environment is improving with time.
@@ III Measurement and Modeling of Detector Noise
To characterize our detectors’ noise over time, we record 2 hours of continuous data following each calibration dataset and remove time periods with abnormally low bias power, which is primarily caused by unusually high environmental vibrations or large energy depositions. Additionally, to select periods without above-threshold events that would contaminate and bias our noise, we find the largest magnitude pulse in the entire 200 ms period using an optimal filter and select periods for which this fit has a negative amplitude, i.e., the largest event in the trace is a statistical fluctuation rather than a real energy deposition. In total, 3.9% of the 200 ms traces pass these very strict noise selection criteria and are used to estimate the noise Cross Power Spectral Density (CSD, see supplementary material section C for further discussion of our cuts). We reference this CSD to noise equivalent power by applying a three-pole responsivity modelLindeman et al. (2011); Irwin and Hilton (2005) for each channel (i.e. $\partial P/\partial I(f)$) that is estimated from IV and $\frac{\partial I}{\partial V}$ measurements as in Ref. Anthony-Petersen et al. (2025).
Similar to our observations in Ref. Anthony-Petersen et al. (2025), we find that the noise is significantly larger than the expected theoretical TES noise Irwin and Hilton (2005) and has a significant broad band channel correlated component (which should be statistically consistent with zero). The broad spectrum correlated noise is consistent with shot noise from many subthreshold phonon events in the substrate; i.e. it has the same frequency shape and channel energy partition ($\sim$50% left, $\sim$50% right) as substrate-interacting calibration photons (Fig. 3 Top Right). The magnitude of this broad correlated noise appears to scale with the detector volume, since the correlated noise is four times larger in the 4 mm detector than in the 1 mm detector (Fig. 3 Top Right).
@@
Figure 3: (Left) Noise in the left channels of the 1 mm (top, green) and 4 mm (bottom, blue) detectors. Our noise is well in excess of our modeled TES noise (gray, dotted), and is composed of correlated phonon noise and uncorrelated noise as in Ref. Anthony-Petersen et al. (2025). The peaks around 150 Hz are due to vibration coupled noise. Grey dashed lines show the primary (electrothermal) pole of the TES. (Top Right) Correlated noise (i.e. off diagonal CSD element $S_{LR}$) in the 4 mm and 1 mm detectors. The correlated noise in the 4 mm detector is 4 times as large as the 1 mm detector. The amplitude of the 1 mm correlated noise is fit in the orange highlighted region (to avoid vibration coupled peaks), while the 4 mm noise is fit in the orange and yellow regions. Dashed green and blue lines show the primary phonon poles of the 1 mm and 4 mm detectors respectively. (Bottom Right) Uncorrelated noise (i.e. $S_{LL}$, with the modeled correlated noise subtracted) in the left channels of the 1 mm and 4 mm detectors, consistent with the modeled TES Johnson noise at high frequencies and an excess noise term that is approximately flat and consistent between the two channels at low frequencies (see section D for further discussion). Data corresponds to the final (day 12) dataset.
@@
We additionally find that the TES bias power (i.e. the power applied to the TES to keep it in transition) increases substantially over time (see Fig. 4). We assume that this increase over time is essentially entirely due to a decreasing parasitic power over time, as shifts in bias power due to a cooling dilution fridge base (mixing chamber, MC) temperature should be very small (see supplementary material section B). As with the correlated noise, the shift in the bias power over time is approximately four times larger in the 4 mm detector than in the 1 mm detector.
As shown in Fig. 4 right, the correlated noise and bias power are found to be linearly related for all times in both detectors and thus share an identical functional dependence on time. This strongly suggests that both observations are caused by the same underlying mechanism.
To gain insight with a simple benchmark model, a random process in which a quantized energy deposition $\varepsilon$ occurs with an average rate $R(t)$ that scales with the detector thickness and $t^{-\kappa}$ will produce a parasitic power and excess shot noise of
@@
Figure 4: Correlated noise level (Top Left) and Left + Right channel bias powers (Bottom Left) in the 1 mm (green) and 4 mm (blue) detectors as a function of time. (Right) Correlated noise level vs. Left + Right bias powers. Dashed lines show fit models, representing shot noise that scales in rate with detector volume with a shot noise quantum of $\varepsilon=0.68\pm 0.38$ meV, comparable to the aluminum superconducting bandgap.
We plot the excess correlated noise and bias power as a function of time and find good agreement with this model (an exponential trend is not a good fit to the data). We find $\varepsilon=0.68\pm 0.38$ meV, and $\kappa=0.635\pm 0.009$. While the characteristic energy scale $\varepsilon$ seems to remain constant over time, there is no reason to assume that the shot noise events all have an identical energy. As discussed in supplementary material section E, $\epsilon$ is readily generalizable to processes which have a distribution of energy depositions, where $\varepsilon=\frac{<E^{2}>}{<E>}$. So long as the distribution is not e.g. double peaked, $\varepsilon$ can still be thought of as a characteristic energy scale of the underlying phonon burst process.
[…]
Figure 5: (Top) The mass normalized shared background rates in the 1 mm and 4 mm detectors, with backgrounds measured in the SPICE 1%Anthony-Petersen et al. (2025); Chang et al. (2025), CPDv1Fink et al. (2021); Alkhatib et al. (2021), CPDv2 (previously unpublished), and CRESST 0.35g Si Angloher et al. (2023) detectors. Broadly, all six detectors seem to observe the same backgrounds normalizing for mass. Note that as the CPD and CRESST detectors are one channel detectors which cannot reject singlesAnthony-Petersen et al. (2025), the increased background rates near threshold could be due to singles, as well as noise triggers or additional LEE backgrounds. For clarity, the 1 mm, 4 mm, and SPICE 1% detectors have their spectra cut off at 30 eV to remove saturated events from cosmic rays and radioactive backgrounds which begin to appear at these energies. (Top insert) Detail of the 1 mm, 4 mm, and SPICE 1% detector backgrounds. The orange and green bands show the energy ranges in which time dependence was measured (see bottom right). (Left Bottom) Background event energy partitioning between left and right channels, showing single and shared events. Note the energy reconstruction assumes a shared pulse shape in each channel. (Right bottom) Rate of shared events over time in the 1 mm detector in the shaded 2.5-5 eV and 5-10 eV bins in the center top figure. Dashed line shows the (weighted) average rate in the 5-10 eV range, and a power law fit in the 2.5-5 eV bin.
@@ IV Above Threshold Shared Backgrounds
In addition to this phonon burst shot noise from below threshold events, we observe a high rate of above threshold background events (often called the “Low Energy Excess,” or LEEAdari et al. (2022); Baxter et al. (2025)). To investigate these LEE events, we recorded 2 or 12 hour periods of continuous background data following each calibration run, and triggered this dataset offline using an optimum filter optimized to search for events with a channel energy partition and pulse shape consistent with that of substrate absorbed calibration photons as in Ref. Chang et al. (2025). In Fig. 5 bottom left, a scatter plot shows the energy deposited in each channel without a constraint on the partition. Identical to observations in Ref. Anthony-Petersen et al. (2025), the LEE backgrounds split into two classes: “shared” events with an approximately equal channel partition and a pulse shape identical to substrate absorbed photons, and “single” events where nearly all energy is deposited in a single phonon sensor channel. Singles have a pulse shape with a fast rise, but a slow fall consistent with an energy deposition that saturates a subset of the athermal phonon sensors that are read out in parallel within a channel.Anthony-Petersen et al. (2025) Each event is then fit with a multi-channel optimum filter whose pulse shape and channel partition are fixed to the average of each LEE class (singles left, singles right, shared), and classified according to lowest $\Delta\chi^{2}$.Anthony-Petersen et al. (2025); Chang et al. (2025)
Similar to our observations of correlated phonon noise, we see that the rate of above-threshold shared background events scales with the detector thickness, as it does for below threshold phonon shot noise (the y-axis has been normalized by substrate mass, proportional to detector thickness). However, the rate of these high energy above threshold shared events is consistent with being constant with time on these short timescales. At lower energies, where statistics are better, the rate of shared LEE events does appear to decrease over time.
As in Ref. Anthony-Petersen et al. (2025), our singles appear to originate within the aluminum QET films, and we additionally observe that their rate decreases with time (see further discussion in supplementary materials section F). These observations give additional support to the proposal that dislocation mediated relaxation of thermal stress in these aluminum QET films is responsible for these singles. Anthony-Petersen et al. (2025); Romani (2024)
@@ V Discussion
We observe that both the correlated phonon noise level and the shared LEE backgrounds scale with the detector substrate thickness. From this scaling alone, we conclusively reject the hypotheses that our aluminum or tungsten sensor films constitute the dominant source of shared phonon bursts Romani (2024). Likewise, the phonon bursts are unlikely to originate from the top and bottom polished 1 cm2 silicon faces of the substrate.
This thickness scaling observation is compatible with a phonon burst processes whose rate scales with the diced sidewall area, the mass, or the volume of the substrate. To discriminate between these hypotheses, we would ideally measure shared LEE and correlated noise rates in detectors with different volume to sidewall ratios, while keeping all other detector characteristics fixed. Unfortunately, this study with strict extraneous variable control has not been done.
However, we have characterized the LEE rate in two large area Cryogenic Photon Detectors (CPDs) Alkhatib et al. (2021); Fink et al. (2021). These large area cylindrical substrates (diameter: 76.2 mm, thickness 1 mm) have $\sim\times 8$ larger volume to sidewall ratio than the 1 cm2 detectors discussed to this point. We observe good agreement with volume scaling, and find that when scaling by sidewall area, both noise and above threshold backgrounds differ between these two sets of detectors by around an order of magnitude. While the differences in holding method, fabrication, handling, and storage history detract from our ability to control extraneous variables, we believe that the evidence for volume scaling is compelling (see the supplementary materials H for additional discussion of these systematics).
@@
Volume scaling suggests that defects in the detector substrate are responsible for LEE and excess phonon noise observations. For example, the relaxation of neutron induced crystal damage discussed in Refs. Nordlund et al. (2024); Jay et al. (2017) could create phonon bursts consistent with our observations. The strong rate vs. time since cooldown dependence we observe in our noise data, for example, could be explained by assuming that meV-scale gaps between different energy levels in defects are thermally populated at 300 K, and then relax to lower energy states at mK temperatures. The observation of time dependence in above threshold LEE rates, both here and previouslyAngloher et al. (2023), is more difficult to explain given their eV scale energies are much larger than room temperature energy scales. As discussed in Ref. Nordlund et al. (2024), an avalanche-like process where the relaxation of a meV-scale state triggers the release of many more such states is one possible mechanism by which the relaxation of thermally populated states can trigger eV scale events.
Since the shared LEE phonon bursts we observe have been shown to originate in the silicon substrate, it is likely that they would also occur in other devices constructed on silicon substrates such as superconducting qubits and MKID detectors. Phonon bursts produced in the silicon substrate would be expected to create quasiparticles in the device’s superconductors, contributing to the excess equilibrium quasiparticle density long observed in these devices. To estimate the importance of this quasiparticle generation mechanism compared to other sources, we assume that the reduced quasiparticle density $x_{qp}=n_{qp}/n_{cp}=n_{qp}/(4\times 10^{6}\mu\mathrm{m}^{-3})$Martinis (2021) in an aluminum superconductor follows the dynamical equation
$\frac{dx_{qp}}{dt}=-rx_{qp}^{2}+g$ (5)
where $r\approx(20\mathrm{ns})^{-1}$ is the quasiparticle recombination rate and $g$ is the normalized quasiparticle generation rate Martinis (2021); Vepsäläinen et al. (2020); Wang et al. (2014). We assume that the quasiparticle dynamics in the superconductor are dominated by recombination as in Ref. Vepsäläinen et al. (2020).
Modeling the device as a silicon substrate with an aluminum superconducting region (including qubits, KIDs and ground planes) on the surface of the chip, we assume for simplicity that LEE events are evenly distributed through the bulk silicon chip and collected uniformly in the aluminum. This produces an average normalized quasiparticle generation rate of
$g>=\frac{<\rho_{p}>}{2\Delta_{Al}n_{cp}}\frac{V_{Si}}{V_{Al}}$ (6)
@@
$x_{qp}>=\sqrt{\frac{<g>}{r}}=\sqrt{\frac{<\rho_{p}>}{2\Delta_{Al}n_{cp}}\frac{V_{Si}}{V_{Al}}\frac{1}{r}}$ (7)
For the aluminum KID on a silicon substrate described in Ref. Temples et al. (2024), we expect a residual density of $x_{qp}\approx 4\times 10^{-6}$ using this estimate, which compares favorably to the $x_{qp}=4.6\times 10^{-6}$ that they observed in their KID resonator. This technique can also be used to estimate the residual quasiparticle density caused in qubits (taking into account both the $\mathcal{O}$(100 nm) thick qubit and ground plane, which cover nearly all the surface of the chip), and gives $x_{qp}\approx 10^{-8}-10^{-7}$, roughly in line with observations for modern non-gap-engineered qubitsSerniak et al. (2018). Our observations also agree with the phonon origin Yelton et al. (2025) and reduction in rate over time Mannila et al. (2022); Yelton et al. (2025) previously observed for such bursts in superconducting quantum devices constructed on silicon substrates.